Constants Part A A stockroom worker pushes a box with mass 11.3 kg on a horizontal surface with a constant speed of 3.40 m/s. The coeflficient of kinetic friction between the box and the surface is 0.19 What horizontal force must the worker apply to maintain the motion? Express your answer with the appropriate units. |F-1 Value Units Submit Request Answer
Answer #1

Apply Newton second law to the box

F- fk = ma

since box moves with constant speed a = 0

F- fk = 0

F = fk

= uk N

= uk mg

= 0.19( 11.3) (9.8)

= 21.04 N